For a given company, both this year's and next year's NOPAT is anticipated to be $1000, and it has current invested capital of $5,000. The company expects annual growth going forward of 5%, and its weighted-average cost of capital is 13%. Using the Key Driver Formula (KDF), what is the value of this company?

Round your answer to the nearest whole number and do not include a dollar sign or comma(s).

With this information about the company from the previous question, what is that company's economic profit?
Answer #1

NOPAT = 1000
Growth = 5%
WACC = 13%
Return on Net invested Capital (RONIC) = 1000/5000 = 20%

Value of this company = NOPAT*(1-Growth/RONIC)/(WACC-Growth) = 1000*(1-5%/20%)/(13%-5%) = 9375

Economic Profit = (RONIC-WACC)*Capital invested = (20%-13%)*5000 = 350

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well